OpenStudy (anonymous):

Circle c has a radius of 10cm. Each of pounts B and D is on the midpoint of the radius. Find the area of the shaded region. Please how come the answer is 45.19? Can someone give me the solution for this? Please please thank youuu!

OpenStudy (accessdenied):

Can you show us the shaded region in the question?

OpenStudy (anonymous):

yes, we can't help you if you don't show us a picture of the shaded area

OpenStudy (anonymous):

Please help me :( i can't find the area of the small triangle :( please show me the solution why the answer is 45.19 :( thanks guys

OpenStudy (anonymous):

please help me :(

OpenStudy (accessdenied):

By small triangle, you mean this one?  And can you find the area of the kite figure underneath it?

OpenStudy (anonymous):

 area of the whole quarter of a circle= A=1/4pi (10cm)^2 = 78.5 area of ACB = 10*5/2 = 25 area of BEF = (5* 10/3) / 2 + 8.3333 area of shaded= 79.5 - 25 - 8.333 = 45.1

OpenStudy (anonymous):

why did you divide it by 3? in the area of BEF and where does 8.333 came from? :))

OpenStudy (anonymous):

why did you divide it by 3? and where does 8.333 came from? please master answer me :)

OpenStudy (anonymous):

i found the intersection point between AB and DE. that y-coordinate is the height of the BEF triangle. the base in BEF is also 5. the y-coordinate i found (=the height) is 10/3. so the area A= (base)(height)/2 = (5)(10/3) / 2 = 8.3333